what is the solution of the following system of linear equations?
2x+3y=5
-4x-6y=6

Answer by checkley75(3666)

2x+3y=5
3y=-2x+5
y=-2x/3+5/3
-------------------
-4x-6y=6
-6y=4x+6
y=4x/-6+6/6
y=-2x/3+1
-------------------
(graph 300x200 pixels, x from -6 to 5, y from -10 to 10, of TWO functions y = -2x/3 +5/3 and y = -2x/3 +1).